# Biologically-Inspired Technologies: Integrating Brain-Computer Interface and Neuromorphic Computing for Human Digital Twins

### v1

@@ Abstract
The integration of the Metaverse into a human-centric ecosystem has intensified the need for sophisticated Human Digital Twins (HDTs) that are driven by the multifaceted human data. However, the effective construction of HDTs faces significant challenges due to the heterogeneity of data collection devices, the high energy demands associated with processing intricate data, and concerns over the privacy of sensitive information. This work introduces a novel biologically-inspired (bio-inspired) HDT framework that leverages Brain-Computer Interface (BCI) sensor technology to capture brain signals as the data source for constructing HDT. By collecting and analyzing these signals, the framework not only minimizes device heterogeneity and enhances data collection efficiency, but also provides richer and more nuanced physiological and psychological data for constructing personalized HDTs. To this end, we further propose a bio-inspired neuromorphic computing learning model based on the Spiking Neural Network (SNN). This model utilizes discrete neural spikes to emulate the way of human brain processes information, thereby enhancing the system’s ability to process data effectively while reducing energy consumption. Additionally, we integrate a Federated Learning (FL) strategy within the model to strengthen data privacy. We then conduct a case study to demonstrate the performance of our proposed twofold bio-inspired scheme. Finally, we present several challenges and promising directions for future research of HDTs driven by bio-inspired technologies.
@@
Human Digital Twin, Human-Centric Metaverse, Brain-Computer Interface, Spiking Neural Networks.
@@ I Introduction
The Metaverse, once a concept confined to science fiction, has become a tangible reality. This shared virtual environment, created by the convergence of digitally enhanced physical reality, augmented reality (AR), and ubiquitous Internet access, allows users to engage in interactive, work-related, and social activities within three-dimensional spaces that transcend physical and geographic boundaries. With the advent of Beyond 5G (B5G) and 6G technologies, the Metaverse is evolving into a more human-centric space, i.e., human-centric Metaverse. This new paradigm aims not only to replicate but also to enhance human interactions through sophisticated digital extensions, ensuring that personal and collective engagements within these virtual spaces are as nuanced and meaningful as those in the physical world. To this end, the human digital twins (HDTs) has been proposed to thoroughly bridge the virtual and physical worlds [1].
@@
Different from the virtual representation of users in the Metaverse (i.e., avatars), the HDT is an advanced system that extends beyond mirroring physical appearances. This system integrates and analyzes comprehensive data types of user, including behavioral, psychological, physiological, and observational information, enabling the HDT to adapt dynamically not only to user interactions and immersion within the Metaverse but also to provide real-time monitoring of human behavior and states. By leveraging each user’s unique and real-time data, HDTs efficiently capture real-time feedback and delivers personalized services, offering tailored solutions that meet the needs of both individual users and Metaverse Service Providers (MSPs). Therefore, this system not only significant elevates the Quality of Experience (QoE) perceived by the users but also assists in optimizing the Quality of Service (QoS) provided by the MSPs, ensuring efficient system performance. These multiple benefits highlight the integral role of HDTs in enhancing the overall functionality and sustainability of the human-centric Metaverse.
@@
To construct an effective HDT, users must deploy multiple Internet-of-Things (IoT) devices (e.g., smartwatches, smartphones, and body sensors) to continuously collect real-time data on their physiological and behavioral states. This data is then transmitted to an MSP for comprehensive analysis, enabling the creation of highly personalized HDTs. However, this process presents significant challenges in two main stages respectively, i.e., data collection and data processing. Regarding data collection, the IoT devices used in this process are typically produced by different manufacturers, each operating on distinct software platforms and employing. This diversity complicates the efficient integration and processing of data. Varied operational protocols and data formats can lead to inconsistencies and gaps in the collected data, thereby reducing the efficacy of data collection. Additionally, the sensors in these devices are typically designed to detect only specific physiological signals, such as heart rate and blood oxygen levels. However, this may be insufficient in the dynamic real-world environments where human behaviors and states, such as psychological states like emotions and physiological states like fatigue, are complex and variable. Such sensors have limitations in capturing nuanced psychological or emotional states, which are essential for the immersive human-centric ecosystem.
@@ II Human Digital Twin
In this section, we introduce the human-centric Metaverse and review the current HDT frameworks and their applications. Then, we highlight the key challenges that must be addressed for them to realize their full potential within the human-centric Metaverse.
@@ II-A Human-centric Metaverse and Human Digital Twin Framework
The current HDT framework operates cyclically by executing the following steps: (1) Sensors embedded in IoT devices continuously collect user real-time behavioral data; (2) The data are then transmitted to the MSP; (3) The MSP processes and analyzes the data using advanced software and algorithms. After analyzing these data, MSPs are able to generate the user’s HDT. Currently, this framework has been utilized primarily in healthcare and gaming. For example, HDTs are integrated with wearable devices such as smartwatches, fitness trackers, and health sensors to continuously monitor vital factors like heart rate, blood pressure, and glucose levels. This data can be synchronized with medical records and utilized to generate real-time insights into an individual’s health, enabling early diagnosis of potential issues and the optimization of personalized healthcare strategies [1, 3, 4].
@@ II-B Challenges of the Current Human Digital Twins
While the concept of HDTs offers immense advantages, it is still in its early stages. Many ongoing multifaceted research challenges need to be addressed to fully realize its potential.
@@ II-B1 Devices and Data Sources (Q1)
One of the significant challenges facing HDTs is the heterogeneity of the IoT devices used to collect human data. These IoT devices such as smartwatches, body sensors, smartphones, and fitness trackers are critical for collecting continuous streams of physiological data. However, these devices often produced by different manufacturers and have varying technical specifications. They operate on diverse hardware platforms, utilize different communication protocols, and employ distinct software architectures, creating significant barriers to seamless data integration and real-time processing. This heterogeneity complicates the creation of accurate, responsive HDTs, as each device requires specific processing capabilities and tailored integration efforts. Moreover, the need for real-time synchronization of data from multiple sources introduces additional technical complexities, including latency, bandwidth constraints, energy, and data loss, all of which degrade system performance. On the other hand, inconsistent data formats and interoperability issues increase the difficulty for MSPs in developing standardized frameworks for HDT implementation, thereby limiting the scalability and personalization potential of HDTs.
@@ III Brain-Computer Interface-Enhanced Human Digital Twin
In this section, we introduce BCI technology and its current applications across various domains. We then propose a novel Bio-HDT framework that leverages advanced BCI technology, focusing on the potential for improved customization, efficiency, and interaction within the human-centric Metaverse. Finally, we discuss the unique challenges involved in deploying BCIs within HDT frameworks.
@@ III-A Brain Signal and Brain-Computer Interface
  • Entertainment: BCIs are revolutionizing gaming and immersive experiences by enabling users to control avatars and interact with virtual environments using only their thoughts. This technology significantly enhances user immersion and interaction, setting the stage for more intuitive and responsive VR systems [10, 6, 11].
@@ III-B Biologically-inspired Human Digital Twins
Therefore, the Bio-HDT framework not only addresses the challenge of customizing HDTs for individual users (for Q1) but also minimizes energy consumption (for Q2) and reduces the risk of data breaches (for Q3) by eliminating the need for multiple data collection devices. Additionally, the real-time interaction between brain signals and the decision-making systems enhances the accuracy, responsiveness, and immersiveness of HDTs. This new paradigm brings users closer to a fully realized human-centric Metaverse.
III-C Concerns
@@ III-C2 Energy Efficiency (Q2)
III-C3 Data Privacy and Security (Q3)
@@ IV-A Spiking Neural Network
In the context of the Bio-HDT framework and the human-centric Metaverse, SNNs offer significant advantages due to their ability to process information based on the timing and frequency of spikes, closely mirroring the brain’s natural information processing mechanisms. These capabilities enables SNNs particularly adept at interpreting rapidly fluctuating brain signals and recognizing dynamic, time-based patterns that traditional models often overlook. As illustrated in the step (2) of Fig. 1, extracted brain signals are first fed into a brain-like SNN model. This model learns to process and efficiently encode brain activity into multiple spikes, enhancing the real-time responsiveness and accuracy in user states analysis and personalized service delivery. Importantly, the event-driven nature of SNNs, i.e., activating only during specific spike events, significantly reduces energy consumption and computational overhead, enabling SNNs ideally suited for power-sensitive devices such as HMDs and sensors within the Bio-HDT framework (for Q2). More technical details of SNN can be found in  [13].
@@ IV-B SNN-Driven Federated Learning
Specifically, SNN-driven FL can be deployed on local devices, i.e., BCI-driven HMDs. Instead of transmitting raw data to MSPs, each device trains its own SNN model using an FL framework. Note that model training occurs dynamically as users interact with the Metaverse, utilizing idle times or periods of low interaction to process data and refine the model. This strategy minimizes disruptions and maximizes efficiency during periods of active use. Once the local models are sufficiently trained, only the model updates (i.e., gradients) are transmitted to the MSP. There, they are aggregated to improve model performance without exposing the original brain data [14]. By retaining raw brain signals on local devices, the system protects users’ cognitive, emotional, and behavioral data from potential security threats during transmission. This architecture not only ensures a higher level of data privacy but also addresses critical concerns about the vulnerability of sensitive brain data (for Q3).
@@ IV-C2 SNN architecture
We conduct the learning model through SNN as shown in Fig. 2. This model comprises a spiking encoder network and a classifier network. The spiking encoder is designed to mimic neuronal communication, extracting features from brain signals and converting them into firing spikes at various time steps. The identifier network processes these spikes to carry out identification tasks specifically tailored to brain signals. Specifically, efficient identification of brain signals enables MSPs to understand users’ expectations, thus allowing them to provide personalized services.
@@ IV-C3 FL training and baselines
As illustrated in Fig. 4, FL4BCI-SNN demonstrates WQoES improvements of 2.12 and 2.11 times compared to the baseline methods, significantly outperforming both FL4BCI-CNN and FL4BCI-LSTM. This result confirms the superior efficacy of our proposed FL4BCI-SNN model in enhancing WSP. Note that although the constructed WSP metric does not currently integrate with other indicators that influence QoE and QoS, such as resource allocation schemes and wireless communication quality, it is designed to be adaptable and can easily be combined with these factors.
@@ V-A Challenges on Processing Brain Signals
Another critical factor is ensuring real-time synchronization between brain activity and HDTs within the future human-centric Metaverse. This synchronization is crucial for maintaining an immersive and responsive user experience. With anticipated advancements in B5G/6G systems, the Bio-HDT framework is prone to overcome existing latency challenges by enabling near-instantaneous data transmission. The introduction of 6G could transform Bio-HDT interactions, with ultra-low latency, edge computing, and enhanced data rates combining to ensure seamless and immersive experiences in the Metaverse. These technologies will allow for brain signals to be processed and responded to almost instantly, significantly enhancing user interaction.
@@ VI Conclusion
In this work, we have introduced the current development of HDTs and highlighted the gap between existing HDT frameworks and the evolving requirements of the human-centric Metaverse. To bridge this gap, we have proposed a two-layer bio-inspired scheme. First, we introduce BCI as the first-layer of bio-inspired to create the Bio-HDT framework. This framework is able to capture users’ brain signals, which contain underlying information about humans’ multifaceted data. This capability provides richer data and a holistic view for constructing personalized HDTs, while simultaneously reducing the dependency on the diverse data collection devices. To accurately understand user brain activities, we have proposed a neuromorphic computing learning model based on SNN as the second-layer of bio-inspired. This model not only improves the accuracy of brain signals identification but also reduces energy consumption by mimicking the brain’s natural processing methods, thus enhancing the system performance and then enabling the QoE perceived by the users and QoS provided by MSPs. We have then proposed the SNN-driven FL to enhance privacy conservation. Moreover, we have conducted a case study to validate the efficacy of our proposed scheme. Finally, we have discussed some challenges that are worthy of further research. We believe that this work provides valuable insights into the application of bio-inspired techniques in developing personalized Metaverse services, ultimately advancing the construction of a human-centric next-generation Internet.

### v2

@@ Abstract
The integration of immersive communication into a human-centric ecosystem has intensified the demand for sophisticated Human Digital Twins (HDTs) driven by multifaceted human data. However, the effective construction of HDTs faces significant challenges due to the heterogeneity of data collection devices, the high energy demands associated with processing intricate data, and concerns over the privacy of sensitive information. This work introduces a novel biologically-inspired (bio-inspired) HDT framework that leverages Brain-Computer Interface (BCI) sensor technology to capture brain signals as the data source for constructing HDT. By collecting and analyzing these signals, the framework not only minimizes device heterogeneity and enhances data collection efficiency, but also provides richer and more nuanced physiological and psychological data for constructing personalized HDTs. To this end, we further propose a bio-inspired neuromorphic computing learning model based on the Spiking Neural Network (SNN). This model utilizes discrete neural spikes to emulate the way of human brain processes information, thereby enhancing the system’s ability to process data effectively while reducing energy consumption. Additionally, we integrate a Federated Learning (FL) strategy within the model to strengthen data privacy. We then conduct a case study to demonstrate the performance of our proposed twofold bio-inspired scheme. Finally, we present several challenges and promising directions for future research of HDTs driven by bio-inspired technologies.
@@
Human Digital Twin, Human-Centric Systems, Brain-Computer Interface, Spiking Neural Networks.
@@ I Introduction
The paradigm shift in communication networks from 5G to 6G is accelerating the realization of immersive communication. Central to this transformation is the human-centric Metaverse, a shared virtual environment emerging from the seamless fusion of digitally enhanced physical reality, augmented reality (AR), and ubiquitous Internet connectivity. By enabling users to engage in interactive, professional, and social activities within three-dimensional spaces, the Metaverse effectively transcends physical and geographic boundaries. To further bridge the gap between virtual and physical worlds, the concept of human digital twins (HDTs) has emerged. These advanced digital representations not only replicate but also enhance human interactions through sophisticated digital extensions, ensuring that both individual and collective engagements in virtual spaces mirror the depth and nuance of the physical world [1].
@@
Different from the virtual representation of users in the Metaverse (i.e., avatars), the HDT is an advanced system that extends beyond mirroring physical appearances. This system integrates and analyzes comprehensive data types of user, including behavioral, psychological, physiological, and observational information, enabling the HDT to adapt dynamically not only to user interactions and immersion within the Metaverse but also to provide real-time monitoring of human behavior and states. By leveraging each user’s unique and real-time data, HDTs efficiently capture real-time feedback and delivers personalized services, offering tailored solutions that meet the needs of both individual users and Metaverse Service Providers (MSPs). Therefore, this system not only significant enhances the Quality of Experience (QoE) perceived by the users but also assists in optimizing the Quality of Service (QoS) provided by the MSPs, ensuring efficient system performance. These multiple benefits highlight the integral role of HDTs in enhancing the overall functionality and sustainability of the human-centric Metaverse.
@@
To construct an effective HDT, users must deploy multiple Internet-of-Things (IoT) devices (e.g., smartwatches, smartphones, and body sensors) to continuously collect real-time data on their physiological and behavioral states. This data is then transmitted to an MSP for comprehensive analysis, enabling the creation of highly personalized HDTs. However, this process presents significant challenges in two main stages respectively, i.e., data collection and data processing. Regarding data collection, the IoT devices used in this process are typically produced by different manufacturers, each operating on distinct software platforms and utilizing different standards and protocols. This diversity complicates the efficient integration and processing of data. Varied operational protocols and data formats can lead to inconsistencies and gaps in the collected data, thereby reducing the efficacy of data collection. Additionally, the sensors in these devices are typically designed to detect only specific physiological signals, such as heart rate and blood oxygen levels. However, this may be insufficient in the dynamic real-world environments where human behaviors and states, such as psychological states like emotions and physiological states like fatigue, are complex and variable. Such sensors have limitations in capturing nuanced psychological or emotional states, which are essential for the immersive human-centric ecosystem.
@@ II Human Digital Twin
In this section, we introduce the human-centric Metaverse and review the current HDT frameworks and their applications. Then, we highlight the key challenges that must be addressed for realizing their full potential within the human-centric Metaverse.
@@ II-A Human-centric Metaverse and Human Digital Twin Framework
As shown in Fig. 1, the current HDT framework operates cyclically by executing the following steps: (1) Sensors embedded in IoT devices continuously collect user real-time behavioral data; (2) The data are then transmitted to the MSP; (3) The MSP processes and analyzes the data using advanced software and algorithms. After analyzing these data, MSPs are able to generate the user’s HDT. Currently, this framework has been utilized primarily in healthcare. For example, HDTs are integrated with wearable devices such as smartwatches, fitness trackers, and health sensors to continuously monitor vital factors like heart rate, blood pressure, and glucose levels. This data can be synchronized with medical records and utilized to generate real-time insights into an individual’s health, enabling early diagnosis of potential issues and the optimization of personalized healthcare strategies [1, 3, 4].
@@ II-B Challenges of the Current Human Digital Twins
While the concept of HDTs offers immense advantages, it is still in its early stages. Multifaceted challenges need to be addressed to fully realize its potential.
Figure 1: The current HDT framework. Wearable devices collect observational, physiological, and behavioral data. These data are transmitted in real-time to the MSP through wireless links. The MSP then employs a range of customized software and models to construct and continuously update the HDT.
@@ II-B1 Devices and Data Sources (Q1)
One of the significant challenges facing HDTs is the heterogeneity of the IoT devices used to collect human data. These IoT devices such as smartwatches, body sensors, smartphones, and fitness trackers are critical for collecting continuous streams of physiological data. However, these devices are often produced by different manufacturers and have varying technical specifications. They operate on diverse hardware platforms, utilize different communication protocols, and employ distinct software architectures, creating significant barriers to seamless data integration and real-time processing. This heterogeneity complicates the creation of accurate, responsive HDTs, as each device requires specific processing capabilities and tailored integration efforts. Moreover, the need for real-time synchronization of data from multiple sources introduces additional technical complexities, including latency, bandwidth constraints, energy, and data loss, all of which degrade system performance. On the other hand, inconsistent data formats and interoperability issues increase the difficulty for MSPs in developing standardized frameworks for HDT implementation, thereby limiting the scalability and personalization potential of HDTs.
@@ III Brain-Computer Interface-Enhanced Human Digital Twin
In this section, we illustrate BCI technology and its current applications across various domains. We then propose a novel Bio-HDT framework that leverages advanced BCI technology, focusing on the potential for improved customization, efficiency, and interaction within the human-centric Metaverse. Finally, we discuss the unique challenges involved in deploying BCIs within HDT frameworks.
@@ III-A Brain Signal and Brain-Computer Interface
  • Entertainment: BCIs are revolutionizing gaming and immersive experiences by enabling users to control avatars and interact with virtual environments using only their thoughts. This technology significantly enhances user immersion and interaction, setting the stage for more intuitive and responsive VR systems [6, 10].
@@ III-B Biologically-inspired Human Digital Twins
Therefore, the Bio-HDT framework not only addresses the challenge of customizing HDTs for individual users (for Q1) but also minimizes energy consumption (for Q2) and reduces the risk of data breaches (for Q3) by eliminating the need for multiple data collection devices. Additionally, the real-time interaction between brain signals and the decision-making systems enhances the accuracy, responsiveness, and immersiveness of HDTs. This new paradigm brings users closer to a fully realized human-centric Metaverse. It is worth noting that the proposed Bio-HDT system can also serve humans effectively beyond the Metaverse environment. In such cases, the Bio-HDT system can be used to monitor various aspects of human states [7].
III-C Challenges
@@ III-C2 Energy Efficiency
III-C3 Data Privacy and Security
@@ IV-A Spiking Neural Network
In the context of the Bio-HDT framework and the human-centric Metaverse, SNNs offer significant advantages due to their ability to process information based on the timing and frequency of spikes, closely mirroring the brain’s natural information processing mechanisms. These capabilities enables SNNs particularly adept at interpreting rapidly fluctuating brain signals and recognizing dynamic, time-based patterns that traditional models often overlook. As illustrated in the step (2) of Fig. 2, extracted brain signals are first fed into a brain-like SNN model. This model learns to process and efficiently encode brain activity into multiple spikes, enhancing the real-time responsiveness and accuracy in user states analysis and personalized service delivery. Importantly, the event-driven nature of SNNs, i.e., activating only during specific spike events, significantly reduces energy consumption and computational overhead, enabling SNNs ideally suited for power-sensitive devices such as HMDs and sensors within the Bio-HDT framework (for Q2). More technical details of SNN can be found in [12, 13].
@@ IV-B SNN-Driven Federated Learning
Specifically, SNN-driven FL can be deployed on local devices, i.e., BCI-driven HMDs. Instead of transmitting raw data to MSPs, each device trains its own SNN model using an FL framework. Note that the model training is conducted dynamically, leveraging idle periods or times of low interaction to process data and refine the model effectively. This strategy minimizes disruptions and maximizes efficiency during periods of active use. Once the local models are sufficiently trained, only the model updates (i.e., gradients) are transmitted to the MSP. There, they are aggregated to improve model performance without exposing the original brain data [14]. By retaining raw brain signals on local devices, the system protects users’ cognitive, emotional, and behavioral data from potential security threats during transmission. This architecture not only ensures a higher level of data privacy but also addresses critical concerns about the vulnerability of sensitive brain data (for Q3).
@@ IV-C2 SNN architecture
We conduct the learning model based on the SNN framework as shown in Fig. 3. This model comprises a spiking encoder network and a classifier network. The spiking encoder is designed to mimic neuronal communication, extracting features from brain signals and converting them into firing spikes at various time steps. The identifier network processes these spikes to carry out identification tasks specifically tailored to brain signals. Specifically, efficient identification of brain signals enables MSPs to understand users’ expectations, thus allowing them to provide personalized services.
@@ IV-C3 FL training and baselines
As illustrated in Fig. 5, FL4BCI-SNN demonstrates WSP improvements of 2.12 and 2.11 times compared to the baseline methods, significantly outperforming both FL4BCI-CNN and FL4BCI-LSTM. This result confirms the superior efficacy of our proposed FL4BCI-SNN model in enhancing WSP. Note that although the constructed WSP metric does not currently integrate with other indicators that influence QoE and QoS, such as resource allocation schemes and wireless communication quality, it is designed to be adaptable and can easily be combined with these factors.
@@ V-A Challenges on Processing Brain Signals
Another critical factor is ensuring real-time synchronization between brain activity and HDTs within the future human-centric Metaverse. This synchronization is crucial for maintaining an immersive and responsive user experience. With anticipated advancements in B5G/6G systems, the Bio-HDT framework is prone to overcome existing latency challenges by enabling near-instantaneous data transmission. The advent of 6G technology has the potential to revolutionize Bio-HDT interactions by leveraging ultra-low latency, edge computing, and enhanced data rates to deliver seamless and immersive experiences within the Metaverse. These advancements will enable brain signals to be processed and responded to in near real-time, significantly enhancing the depth and responsiveness of user interactions.
@@ VI Conclusion
This paper proposed the bio-inspired scheme to address the challenges in the construction of HDT. Specifically, we have introduced BCI as the data sources to create the Bio-HDT, which provides richer data and a holistic view for constructing personalized HDTs, while simultaneously reducing the dependency on the diverse data collection devices. We have then proposed a neuromorphic computing learning model that improves the accuracy of brain signals identification and reduces energy consumption by mimicking the brain’s natural processing methods, thereby enhancing the system performance. Furthermore, we have proposed the SNN-driven FL to enhance privacy conservation and conducted a case study to validate the efficacy of our proposed scheme. Finally, we have discussed some challenges that are worthy of further research. We believe that this work provides valuable insights into the application of bio-inspired techniques in developing personalized immersive services, ultimately advancing the construction of a human-centric next-generation Internet.